OpenStudy (anonymous):

Find the value of k so that the three points lie on the same line. Write the equation of the line in point-slope form. (1,-2) (-2,4) (4,k)

OpenStudy (anonymous):

k= -8 y+8 = -2 (x - 4)

OpenStudy (anonymous):

How did you come up with k=-8?

OpenStudy (anonymous):

Can someone explain how to solve this?

OpenStudy (anonymous):

Since all three points are on the same line, the slope of the segment connecting any two points will be the same. So, you know that \[\large\underbrace{\frac{4-(-2)}{-2-1}}_{\text{slope between first and second point}}=\underbrace{\frac{k-4}{4-(-2)}}_{\text{slope between second and third point}}\] Solve the equation for \(k\): \[\begin{align*}\frac{4-(-2)}{-2-1}&=\frac{k-4}{4-(-2)}\\ \frac{6}{-3}&=\frac{k-4}{6}\\ -2&=\frac{k-4}{6}\\ -12&=k-4\\ -12+4&=k\\ -8&=k \end{align*}\]

OpenStudy (anonymous):

@JoStudy, to find the equation of the line, you use the point-slope formula: \[y-y_0=m(x-x_0)\] where \(m\) is the slope and \((x_0,y_0)\) is any point that you know is on the line. We've already found out that \(m=-2\) (third line in the work above). Now pick any of the three points for \((x_0,y_0)\), then plug it all into the equation.

OpenStudy (anonymous):

Thanks!